OpenStudy (anonymous):

V= (4/3)pi(r^3), where r is the radius, in centimeters. By approximately how much does the volume of a sphere increase when the radius is increased from 3.0cm to 3.2cm?

OpenStudy (anonymous):

your job is to find \[\frac{3}{4}\pi (3.2)^3-\frac{3}{4}\pi (3)^3\] or \[\frac{3\pi}{4}\left((3.2)^3-27\right)\]

OpenStudy (anonymous):

i would use a calculator

OpenStudy (anonymous):

Where did you get 3/4\[\Pi\] from

OpenStudy (anonymous):

you wrote it

OpenStudy (anonymous):

oh no!!

OpenStudy (anonymous):

i can't read is all.

OpenStudy (anonymous):

\[\frac{4\pi}{3}\left((3.2)^3-27\right)\]

OpenStudy (anonymous):

i have up down dyslexia

OpenStudy (anonymous):

Alright, thank you.